# TF-Restormer: Complex Spectral Prediction for Speech Restoration

### v2

@@ Abstract
Speech restoration in real-world conditions is challenging due to compounded distortions such as clipping, band-pass filtering, digital artifacts, noise, and reverberation, and low sampling rates. Existing systems, including vocoder-based approaches, often sacrifice signal fidelity, while diffusion models remain impractical for streaming. Moreover, most assume a fixed target sampling rate, requiring external resampling that leads to redundant computations. We present TF-Restormer, an encoder-decoder architecture that concentrates analysis on input-bandwidth with a time-frequency dual-path encoder and reconstructs missing high-frequency bands through a light decoder with frequency extension queries. It enables efficient and universal restoration across arbitrary input-output rates without redundant resampling. To support adversarial training across diverse rates, we introduce a shared sampling-frequency-independent (SFI) STFT discriminator. TF-Restormer further supports streaming with a causal time module, and improves robustness under extreme degradations by injecting spectral inductive bias into the frequency module. Finally, we propose a scaled log-spectral loss that stabilizes optimization under severe conditions while emphasizing well-predicted spectral details. As a single model across sampling rates, TF-Restormer consistently outperforms prior systems, achieving balanced gains in signal fidelity and perceptual quality, while its streaming mode maintains competitive performance for real-time use. Anonymous code and demos are available at https://tf-restormer.github.io/demo.
@@ 1 Introduction
Speech enhancement (Ephraim & Malah, 1984; Pascual et al., 2017) has historically progressed through isolated sub-tasks with dedicated models such as denoising (Hu et al., 2020; Ho et al., 2020), dereverberation (Han et al., 2015; Wang & Wang, 2020), declipping (Mack & Habets, 2019), and bandwidth extension or super-resolution (Liu et al., 2022a; Lee & Han, 2021). In real-world settings, however, multiple distortions often coincide and are further compounded by digital distortions including lossy codecs (e.g., MP3, Ogg). These factors obscure magnitude and phase in the signal, making coherent, faithful speech restoration substantially more difficult.
@@
This growing complexity has prompted a shift towards general speech restoration using generative models to handle diverse distortions (Liu et al., 2022b; Serrà et al., 2022). Vocoder (Kumar et al., 2019; Kong et al., 2020a)-based approaches reconstruct waveforms from compressed representations (Liu et al., 2022b; Andreev et al., 2022; Babaev et al., 2024). While such approaches improve perceptual quality, they discard phase information and treat speech as a semantic abstraction rather than a physical signal, producing output signals that deviate significantly from the input. In contrast, waveform-based generative models (Oord et al., 2016; Serrà et al., 2022) including GAN- and diffusion-based methods operate directly on the waveform, avoiding the abstraction bottleneck. Diffusion models (Welker et al., 2022; Richter et al., 2023) directly generate waveforms or spectra with improved fidelity, but their temporal compression or iterative sampling precludes streaming.
@@
A further limitation of existing approaches lies in their assumption of a fixed target sampling rate. Restoration models, as well as dedicated super-resolution methods (Kim et al., 2024; Lu et al., 2025), invariably begin by resampling the input to the target rate before processing. This design choice simplifies model training but introduces critical drawbacks: every input must be converted to the highest rate regardless of its native bandwidth, leading to redundant computation from resampling. Moreover, supporting multiple output rates requires either training separate models or repeatedly upsampling and downsampling, both of which are inefficient and impractical in real-world applications. These limitations call for a universal framework that analyzes input bandwidth directly and flexibly synthesizes arbitrary output rates without external resampling.
@@
On the other hand, for denoising and separation tasks, models have been developed primarily in the complex STFT domain (Choi et al., 2018; Hu et al., 2020). In particular, recent works introduced time-frequency (TF) dual-path models (Dang et al., 2022; Wang et al., 2023), alternating sequence modeling along time and frequency axes. This design has proved effective for enhancement and separation tasks by preserving the frequency structure, and it naturally supports sampling-frequency-independent (SFI) formulations (Paulus & Torcoli, 2022; Zhang et al., 2023), since frequency bins can be treated as sequences whose length scales with the input rate while maintaining a consistent STFT frame duration. However, despite this flexibility, existing models still assume matched input-output rates. Moreover, because TF dual-path models explicitly preserve fine spectral structure, their computational cost grows substantially with higher sampling rates, making it difficult to apply to super-resolution tasks directly.
Figure 1: Overview of TF-Restormer. The model adopts an encoder-decoder design: TF-Encoder performs input-bandwidth analysis through stacked time and frequency modules, while the TF-Decoder reconstructs missing high-frequency bands with lightweight processing and learnable extension queries.
@@
These limitations suggest the need for a new design that retains the strengths of TF dual-path processing while overcoming the inefficiency of fixed input-output rates. To this end, we present TF-Restormer, an encoder-decoder architecture for robust speech restoration under diverse degradations. Inspired by masked autoencoders (MAE) (He et al., 2022), TF-Restormer concentrates heavy processing in a TF dual-path encoder that analyzes the input bandwidth, while a lightweight decoder reconstructs the missing high-frequency components through learnable extension queries (Figure 1) with cross-self attention mechanism (Gupta et al., 2023). This asymmetric design enables arbitrary input-output sampling rates without external resampling, which minimizes redundant computation.
@@
  • We design an asymmetric encoder-decoder framework based on a TF dual-path Transformer with SFI-STFT: the encoder focuses on the input bandwidth, while the lightweight decoder extends high frequencies via learnable queries with a cross-self mechanism. This design enables general speech restoration, including super-resolution across arbitrary input-output rates without external resampling, with a shared SFI-STFT discriminator that supports unified adversarial training across diverse rates.
  • We improve robustness and practicality of TF dual-path processing by enhancing the frequency module with a projection-based spectral inductive bias and extending the time module to a causal variant for streaming, ensuring stable operation under extreme degradations as well as real-time applicability.
  • We propose the scaled log-spectral loss as an auxiliary spectral objective that replaces conventional $\ell_{1}$$\ell_{2}$ spectral objectives and complements perceptual (Babaev et al., 2024) and adversarial training (Mao et al., 2017). By selectively emphasizing reliably predictable regions, this loss stabilizes optimization under severe distortions while mitigating oversmoothing.
@@ 2 Related Work
Vocoder-based restoration
Vocoder systems typically project speech into Mel features (Liu et al., 2022b; Babaev et al., 2024) or learned representations (Koizumi et al., 2023; Li et al., 2024) and synthesize waveforms with neural vocoders (Oord et al., 2016; Kumar et al., 2019; Kong et al., 2020a). They often achieve “studio-like” perceptual quality but reduce fidelity, since intermediate features serve as perceptual cues rather than physical signals. Most also rely on temporally compressed U-Net structures (Pascual et al., 2017; Stoller et al., 2018), limiting real-time use. In contrast, we directly predict complex STFTs within a TF dual-path structure, retaining both magnitude and phase while enabling a streaming variant with minimal changes. Recent studies (Kaneko et al., 2022; Lu et al., 2025) further show that direct spectral prediction with adversarial training can rival vocoder-based generation.
Diffusion-based restoration
Another approach is diffusion-based generation from waveform (Kong et al., 2020b; Serrà et al., 2022; Welker et al., 2022; Scheibler et al., 2024) or STFT inputs (Lemercier et al., 2023; Richter et al., 2023). While effective in perceptual quality, inference remains costly: even fast samplers require multiple denoising steps, hindering real-time use. Moreover, while diffusion excels at sampling diverse modes of a distribution, according to  Babaev et al. (2024), speech enhancement typically seeks the most likely clean realization consistent with the observation, not a diverse set of alternatives. As a result, diffusion introduces unnecessary complexity for restoration or enhancement task. Therefore, we target a main-mode estimate via single-pass complex spectral prediction augmented with an adversarial loss for perceptual sharpness.
@@
For denoising, dereverberation, and separation, TF dual-path models (Dang et al., 2022; Cao et al., 2022) alternate sequence modeling along time and frequency, preserving spectral structure and supporting SFI designs (Zhang et al., 2023). They show strong performance in enhancement (Cao et al., 2022; Lu et al., 2023; Chao et al., 2024) and separation (Wang et al., 2023; Saijo et al., 2024; Shin et al., 2025), where the speech signal remains largely preserved, so prior models used identical block designs for time and frequency without considering domain-specific differences. Under challenging degradations, we address this with a projection-based frequency module that injects spectral inductive bias for more robust high-frequency recovery. In addition, while prior dual-path models assumed matched input-output rates, we extend them with an encoder-decoder structure that supports super-resolution beyond fixed-rate restoration.
@@
Conventional bandwidth extension and super-resolution models (Liu et al., 2022a; Han & Lee, 2022; Kim et al., 2024; Lu et al., 2025) typically assume a fixed target rate, upsampling all inputs before processing to fill missing bands. While effective, this introduces redundant computation and ties each model to a single output rate, requiring repeated downsampling of outputs for different targets. In contrast, our encoder-decoder framework confines heavy processing to the input bandwidth and restores missing high-frequency bands through lightweight extension queries, improving efficiency. Unlike prior work restricted to predetermined rates, our design naturally supports user-specified outputs through the TF dual-path backbone, offering flexibility and enabling training with data from diverse native sampling rates.
@@ 3 TF-Restormer
3.1 SFI Input-Output Formulation
As an SFI model (Paulus & Torcoli, 2022), TF-Restormer addresses arbitrary input sampling rates $f_{E}$ by constructing the STFT with a constant frame duration (SFI-STFT). Unlike conventional SFI that assumes matched input-output rates, we introduce the first decoupled formulation, enabling inference at user-specified output rates $f_{D}$. Concretely, given an input $x\in\mathbb{R}^{1\times N_{E}}$ with sampling rate $f_{E}$, its STFT is $\mathbf{X}\in\mathbb{R}^{F_{E}\times T\times 2}$, where $F_{E}$ and $T$ are the number of frequency bins and frames. TF-Restormer then predicts $\mathbf{Y}\in\mathbb{R}^{F_{D}\times T\times 2}$ corresponding to an output $y\in\mathbb{R}^{1\times N_{D}}$ at sampling rate $f_{D}$, satisfying $f_{E}:f_{D}=(F_{E}-1):(F_{D}-1)$ under the assumption of consistent frame duration.
To ensure universal applicability across sampling rates, we adopt a 40 ms analysis window with a 20 ms hop. This choice is a common unit in speech analysis and, being integer multiples across typical rates, guarantees consistent STFT construction at $\{8,16,22.05,24,32,44.1,48\}$ kHz without requiring resampling. The maximum number of frequency bins $F_{\text{max}}$ is 961 for $f_{E}=48$kHz.
Figure 2: Overall architecture of TF-Restormer. With SFI-STFT and iSTFT, arbitrary input-output sample rates can be addressed in a single model. The encoder feature serves as both the input and the fixed key/value for the Freq. cross-self module in the TF-Decoder.
@@ 3.2 Analysis Encoder and Extension Decoder
As illustrated in Figure 1 and 2, the TF-Restormer is constructed with TF-encoder and TF-decoder. The TF-encoder is responsible for analyzing the speech component from the input signals $\mathbf{X}\in\mathbb{R}^{F_{E}\times T\times 2}$. Before the TF-encoder, the input complex representation $\mathbf{X}\in\mathbb{R}^{F_{E}\times T\times 2}$ is first projected to $C_{E}$ dimension by 2d convolution (Conv2D) layer with kernel size of (3,3), followed by layer normalization (LN) (Ba et al., 2016). Then, sinusoidal positional embeddings are added along the frequency axis (Freq. PE). In the encoder, the projected feature is alternately processed by freq and time modules $B_{E}$ times to capture speech component.
@@
Then, the encoder features $\mathbf{Z}\hskip-2.84526pt\in\hskip-2.84526pt\mathbb{R}^{F_{\hskip-0.85358ptE}\times T\times C_{\hskip-0.85358ptE}}$ are embedded to both the input of decoder by projection layer and key/value for cross-attention in the frequency cross-self module as shown in Figure 2. Then, the projected encoder features are padded with learnable extension query $\mathbf{q}_{\text{ext}}\hskip-2.84526pt\in\hskip-2.84526pt\mathbb{R}^{(F_{D}-F_{E})\hskip-0.85358pt\times\hskip-0.85358ptT\hskip-0.85358pt\times\hskip-0.85358ptC_{D}}$ whose values are shared across all the frames and processed by frequency and time module by $B_{D}$ times. In particular, the frequency module performs cross-attention based on encoder feature as key/value. Then, the complex STFT values $\mathbf{Y}\in\mathbb{R}^{F_{D}\hskip-0.85358pt\times\hskip-0.85358ptT\hskip-0.85358pt\times\hskip-0.85358pt2}$ are estimated from decoder features by Conv2D layer.
@@ 3.3 Asymmetric TF dual-path module
In TF dual-path modules, given the feature with shape of $\mathbb{R}^{T\hskip-0.56905pt\times\hskip-0.56905ptF\hskip-0.56905pt\times\hskip-0.56905ptC}$ where $F\hskip-2.84526pt\in\hskip-2.84526pt\{F_{E},F_{D}\}$, time modules process $F$ independent sequences with lengths of $T$ while frequency modules consider the feature as $T$ sequences with lengths of $F$ as illustrated in Figure 1. However, frequency bins are usually more static and structured sequence with a fixed length compared to time frames, exhibiting a relatively consistent structural roles. Therefore, we introduce a frequency module with projection-based inductive bias, explicitly encoding structural regularities across frequency bins for stable restoration.
Figure 3: Unit modules in TF-Encoder and TF-Decoder. The (a) time module is based on MHSA with RoPE while (b) the frequency encoder module is based on MHSA with frequency projection layer. (c) The frequency decoder module employs MHCA based on key/value from the encoder features
We adopt common structure for both time and frequency modules borrowing from TF-Locoformer (Saijo et al., 2024). As shown in Figure 3, it consists of two macaron-style (Lu* et al., 2019) convolution feed-forward network (ConvFFN) with Conv1D with kernel size of $K$ for capturing local contexts. In ConvFFN, the expansion factor is 3 with SwiGLU as hidden activation. Between ConvFFN modules, multi-head self-attention (MHSA) is used for global contexts with $H$ heads. The time module performs MHSA on temporal frames with rotary positional encoding (RoPE) (Su et al., 2024) to offer the relative positions while the frequency module applies MHSA on frequency bins with the frequency projection layer to induce the structural bias.
@@
For frequency cross-self module, we replace the first F-ConvFFN in the frequency self module with multi-head cross-attention (MHCA) based on key-value from the encoder feature $\mathbf{Z}_{\text{enc}}$ while query is high-frequency padded region by $\mathbf{q}_{\text{ext}}$, inspired by cross-self attention (Gupta et al., 2023). Therefore, MHCA conditionally operates when $f_{E}<f_{D}$ and extension query is padded, otherwise, bypassed.
@@
Linformer (Wang et al., 2020) introduced linear projections of key-value to reduce the computations of attention, while MLP-Mixer (Tolstikhin et al., 2021) went further by replacing MHSA with static linear operations. Motivated by these insights, we incorporate a frequency linear projection (Fig. 3(c)) to impose an inductive bias for the structural consistency of frequency bins on top of the benefits of dynamic attention. Since frequency bins exhibit consistent characteristics, we share the same projection layer across all modules and key-value mappings. Formally, for each head $h=1,\dots,H$, given key-value $\mathbf{K}_{h,c},\mathbf{V}_{h,c}\in\mathbb{R}^{T\times F}$ at channel $c$, a learnable projection matrix $\mathbf{A}_{h}\hskip-2.27621pt\in\hskip-1.42262pt\mathbb{R}^{{F_{\text{max}}}\times F_{\text{proj}}}$ with dimension $F_{\text{proj}}$ and maximum bins $F_{\text{max}}$ is applied as
$\tilde{\mathbf{K}}_{h,c}$ $=$ $[\mathbf{K}_{h,c},\mathbf{O}]\mathbf{A}_{h}\in\mathbb{R}^{T\times{F}_{\text{proj}}},\quad 1\leq c\leq C_{h},$ (1)
$\tilde{\mathbf{V}}_{h,c}$ $=$ $[\mathbf{V}_{h,c},\mathbf{O}]\mathbf{A}_{h}\in\mathbb{R}^{T\times{F}_{\text{proj}}},\quad 1\leq c\leq C_{h},$ (2)
@@
The modular design of TF dual-path model further enables a seamless extension to streaming mode by replacing the time module with Mamba (Gu & Dao, 2024) blocks. Refer to Appendix C for detailed model configurations.

### v3

# Query-Based Asymmetric Modeling with Decoupled Input-Output Rates for Speech Restoration

@@ Abstract
Speech restoration in real-world conditions is challenging due to compounded distortions and mismatches between input and desired output rates. Most existing systems assume a fixed and shared input–output rate, relying on external resampling that confines learning to a specific target resolution and limits model generality. We address this setting by formulating speech restoration under decoupled input–output rates, and propose TF-Restormer, a query-based asymmetric modeling framework. The encoder concentrates analysis on the observed input bandwidth using a time–frequency dual-path architecture, while a lightweight decoder reconstructs missing spectral content via frequency extension queries. This design enables a single model to operate consistently across arbitrary input–output rate pairs without redundant resampling. Experiments across diverse sampling rates, degradations, and operating modes show that TF-Restormer maintains stable restoration behavior and balanced perceptual quality, including in real-time streaming scenarios. Anonymous code and demos are available at https://tf-restormer.github.io/demo.
Machine Learning, ICML
@@ 1 Introduction
Speech enhancement (Ephraim and Malah, 1984; Pascual et al., 2017) has historically progressed through isolated sub-tasks with dedicated models such as denoising (Hu et al., 2020; Ho et al., 2020), dereverberation (Han et al., 2015; Wang and Wang, 2020), declipping (Mack and Habets, 2019), and bandwidth extension or super-resolution (Liu et al., 2022a; Lee and Han, 2021). In real-world settings, multiple distortions often coincide and further obscure both magnitude and phase, making coherent and faithful speech restoration substantially more difficult. Crucially, this complexity is often accompanied by a fundamental mismatch between the native input bandwidth and the desired output sampling rate, a challenge that remains largely unaddressed by conventional task-specific frameworks.
@@
This growing complexity has prompted a shift towards general speech restoration using generative models to handle diverse distortions (Liu et al., 2022b; Serrà et al., 2022). Vocoder (Kumar et al., 2019; Kong et al., 2020a)-based approaches reconstruct waveforms from compressed representations (Liu et al., 2022b; Babaev et al., 2024). While such approaches improve perceptual quality, they discard phase information and treat speech as a semantic abstraction rather than a physical signal. In contrast, waveform or complex-spectrum-based generative models (Oord et al., 2016; Serrà et al., 2022; Welker et al., 2022; Richter et al., 2023) including GAN- and diffusion-based methods operate directly on the waveform or complex STFT with improved fidelity, but their temporal compression or iterative sampling precludes streaming.
@@
Beyond the choice of generative modeling paradigm, a critical and often overlooked limitation of existing approaches lies in their assumption of a fixed and shared input–output sampling rate. Restoration models, as well as dedicated super-resolution methods (Kim et al., 2024; Lu et al., 2025), typically begin by resampling the input signal to a predetermined target rate before processing. While this design choice simplifies training by fixing the spectral resolution, it also tightly couples the learned representations to a specific output grid. As a result, every input must be converted to the target rate regardless of its native bandwidth, leading to redundant resampling and resolution-specific modeling. Moreover, supporting multiple output rates under this paradigm requires either training separate models or repeatedly upsampling and downsampling, further entangling the learning process with a fixed resolution assumption. These limitations call for a universal framework that analyzes the observed input bandwidth directly and synthesizes arbitrary output rates without relying on external resampling.
@@
In parallel, denoising and separation models have been extensively developed in the complex STFT domain (Choi et al., 2018; Hu et al., 2020). More recently, time–frequency (TF) dual-path models have been introduced (Dang et al., 2022; Wang et al., 2023), which alternate sequence modeling along the time and frequency axes. By treating frequency bins as sequences rather than static feature dimensions, these models preserve spectral structures and achieve strong performance. This formulation naturally supports sampling-frequency-independent (SFI) processing (Paulus and Torcoli, 2022; Zhang et al., 2023), as the frequency axis can be viewed as a sequence scaling with the input sampling rate while maintaining a consistent STFT frame duration. However, existing TF models still assume matched input–output rates. Moreover, because they preserve fine-grained spectral representations throughout the network, their computational cost grows substantially at higher sampling rates, limiting direct application to super-resolution.
Figure 1: Overall architecture of TF-Restormer. The framework employs a query-based asymmetric design to handle decoupled input–output rates. The heavy TF-Encoder focuses on analysis within the native input bandwidth, while the lightweight TF-Decoder reconstructs missing high-frequency bands using learnable extension queries, bypassing the need for redundant resampling.
@@
These observations motivate a new design that retains the strengths of TF dual-path processing while addressing a setting in which the standard learning assumption of matched input–output resolutions no longer holds. To this end, we propose TF-Restormer, an asymmetric encoder–decoder architecture for speech restoration under diverse degradations. Inspired by masked autoencoders (MAE) (He et al., 2022), TF-Restormer concentrates heavy computation in a TF dual-path encoder that analyzes the observed input bandwidth, while a lightweight decoder reconstructs missing high-frequency components through learnable extension queries using a cross-self attention mechanism (Gupta et al., 2023). By explicitly separating input-band analysis from output-band synthesis, this asymmetric design enables a single model to operate across arbitrary input–output sampling rate pairs without external resampling, while preserving physical signal fidelity.
@@
  • We formulate speech restoration under decoupled input–output sampling rates, where the standard learning assumption of matched spectral resolutions no longer holds, and propose a query-based asymmetric TF encoder–decoder that analyzes the input bandwidth and synthesizes arbitrary output rates without external resampling.
  • Our asymmetric design enables a single model to support general speech restoration including super-resolution, across heterogeneous input–output rate pairs, with unified training facilitated by a shared sampling-frequency-independent (SFI) STFT discriminator.
  • We improve robustness and practicality by enhancing the frequency module with a projection-based spectral inductive bias and extending the time module to a causal variant, enabling stable restoration under severe degradations and real-time streaming inference.
  • We propose a scaled log-spectral loss as an auxiliary objective that stabilizes training under extreme distortions while mitigating oversmoothing, complementing perceptual (Babaev et al., 2024) and adversarial training (Mao et al., 2017).
@@ 2 Related Work
Vocoder- and diffusion-based restoration
Generative approaches have significantly improved speech quality. Vocoder systems typically project speech into Mel features (Liu et al., 2022b; Babaev et al., 2024) or learned representations (Koizumi et al., 2023; Li et al., 2024) and synthesize waveforms with neural vocoders (Oord et al., 2016; Kumar et al., 2019; Kong et al., 2020a). They often achieve “studio-like” perceptual quality but reduce signal fidelity, since intermediate features serve as perceptual cues rather than physical signals. Diffusion-based methods from waveform (Kong et al., 2020b; Serrà et al., 2022; Welker et al., 2022; Scheibler et al., 2024) or STFT inputs (Lemercier et al., 2023; Richter et al., 2023) excel at generating diverse realizations. However, each approach often relies on temporal compression or iterative sampling, limiting real-time streaming. More critically, both paradigms are typically bound to a fixed-rate assumption, requiring external resampling to match a specific target bandwidth. In contrast, we propose a single-pass complex spectral prediction framework that avoids the abstraction bottleneck of vocoders and the complexity of diffusion, while supporting variable input–output rates through an asymmetric design.
@@
TF dual-path models (Dang et al., 2022) alternate sequence modeling along time and frequency, preserving spectral structure and naturally supporting sampling-frequency-independent (SFI) designs (Zhang et al., 2023). They have shown strong performance in enhancement (Cao et al., 2022; Lu et al., 2023; Chao et al., 2024) and separation (Wang et al., 2023; Saijo et al., 2024; Shin et al., 2025). However, most prior models employ symmetric block designs for time and frequency processing, despite the distinct statistical characteristics of the two domains. Under challenging degradations, this limitation becomes more pronounced, motivating the use of projection-based frequency modules that inject spectral inductive bias for more robust high-frequency recovery. Furthermore, while existing dual-path models assume matched input–output rates, we extend this paradigm with an asymmetric encoder–decoder structure that explicitly supports restoration under decoupled rates.
@@
Conventional super-resolution models (Liu et al., 2022a; Han and Lee, 2022; Kim et al., 2024; Lu et al., 2025) typically assume a fixed target rate, upsampling inputs before processing to fill missing bands. While effective, this introduces redundancy and ties each model to a single output rate. Our framework moves beyond this fixed-rate constraint by framing super-resolution as a conditional spectral completion. Using frequency extension queries within an asymmetric encoder–decoder, TF-Restormer confines heavy processing to the input bandwidth and restores missing bands as needed, supporting arbitrary, user-specified output rates without rate-specific models.
@@ 3 TF-Restormer
3.1 SFI input–output formulation
As an SFI model (Paulus and Torcoli, 2022), TF-Restormer addresses arbitrary input sampling rates $f_{E}$ by constructing STFT with a constant frame duration (SFI-STFT). Unlike conventional SFI that assumes matched input–output rates, we introduce the first decoupled formulation, enabling inference at user-specified output rates $f_{D}$. Given an input $x\hskip-1.42262pt\in\hskip-1.42262pt\mathbb{R}^{1\times N_{E}}$ with sampling rate $f_{E}$, its STFT is $\mathbf{X}\hskip-1.42262pt\in\hskip-1.42262pt\mathbb{R}^{F_{E}\times T\times 2}$, where $F_{E}$ and $T$ are the number of frequency bins and frames. TF-Restormer then predicts $\mathbf{Y}\hskip-1.42262pt\in\hskip-1.42262pt\mathbb{R}^{F_{D}\times T\times 2}$ corresponding to an output $y\hskip-1.42262pt\in\hskip-1.42262pt\mathbb{R}^{1\times N_{D}}$ at sampling rate $f_{D}$, satisfying $f_{E}\hskip-1.42262pt:\hskip-1.42262ptf_{D}\hskip-1.42262pt=\hskip-1.42262pt(F_{E}\hskip-1.42262pt-\hskip-1.42262pt1)\hskip-1.42262pt:\hskip-1.42262pt(F_{D}\hskip-1.42262pt-\hskip-1.42262pt1)$ under the assumption of consistent frame duration. To ensure universal applicability across sampling rates, we adopt a 40 ms analysis window with a 20 ms hop. This choice is a common unit in speech analysis and, being integer multiples across typical rates, guarantees consistent STFT construction at $\{8,16,22.05,24,32,44.1,48\}$ kHz without requiring resampling. The maximum number of frequency bins $F_{\text{max}}$ is 961 for $f_{E}=48$kHz.
@@ 3.2 Analysis encoder and extension decoder
As illustrated in Figure 1, the TF-Restormer is constructed with TF-encoder and TF-decoder. The TF-encoder is responsible for analyzing the speech component from the input signals $\mathbf{X}$. Before the TF-encoder, the input complex representation $\mathbf{X}$ is first projected to $C_{E}$ dimension by 2d convolution (Conv2D) layer with kernel size of (3,3), followed by layer normalization (LN) (Ba et al., 2016). Then, positional embeddings are added along the frequency axis. In the encoder, the projected feature is alternately processed by freq. and time self modules $B_{E}$ times to capture speech component.
@@
The encoder features $\mathbf{Z}\in\mathbb{R}^{F_{E}\times T\times C_{E}}$ are projected into the decoder space with $C_{D}$ to serve as both the initial features and the key/value pairs for the frequency cross-self module. To reconstruct the missing high-frequency regions beyond the observed bandwidth, we introduce learnable extension queries, $\mathbf{Q}_{\text{ext}}=[\mathbf{q}_{F_{E}+1},\dots,\mathbf{q}_{F_{D}}]^{T}\in\mathbb{R}^{(F_{D}-F_{E})\times C_{D}}$, where $\mathbf{q}_{f}\in\mathbb{R}^{C_{D}}$ is a frequency-wise learnable vector. Instead of simple zero-padding, these queries function as frequency-specific latent priors optimized to capture the characteristics of each sub-band. By slicing these queries from a unified master vector $\tilde{\mathbf{Q}}_{\text{ext}}$ defined over $F_{\text{min}},F_{\text{max}}]$, the decoder maintains consistent spectral alignment regardless of the specific input or output sampling rates. This ensures that the model treats a specific frequency bin with the same learned representation, facilitating rate-invariant generalization. In the frequency cross-self module, these queries explicitly guide the cross-attention mechanism to retrieve relevant acoustic context from the encoder features. Finally, the complex STFT values $\mathbf{Y}$ are estimated from the decoder features through a Conv2D layer.
Figure 2: Unit modules in TF-encoder and decoder is based on (a) the time self module based on MHSA with RoPE and (b) the frequency module based on MHSA with frequency projection layer. The frequency cross-self module employs MHCA based on key/value from the encoder features while frequency self module is based on two ConvFFNs similarity to time self module.
@@ 3.3 Asymmetric TF dual-path module
In TF dual-path modules, given the feature with shape $\mathbb{R}^{T\hskip-0.56905pt\times\hskip-0.56905ptF\hskip-0.56905pt\times\hskip-0.56905ptC}$ where $F\hskip-2.84526pt\in\hskip-2.84526pt\{F_{E},F_{D}\}$, time modules process $F$ sequences with lengths of $T$ while frequency modules consider the feature as $T$ sequences with lengths of $F$ as illustrated in Figure 1. As shown in Figure 2(a), time self module consists of two macaron-style (Lu* et al., 2019) convolution feed-forward network (ConvFFN) with Conv1D with kernel size $K$ for capturing local contexts (Saijo et al., 2024). In ConvFFN, the expansion factor is 3 with SwiGLU as hidden activation. Between ConvFFN modules, multi-head self-attention (MHSA) is used for global contexts with $H$ heads. The time module performs MHSA on temporal frames with rotary positional encoding (RoPE) (Su et al., 2024) to offer the relative positions. On the other hand, the frequency self module applies MHSA with the frequency projection layer to induce the structural bias as frequency bins are more static sequence with a fixed length, exhibiting a relatively consistent structural roles.
@@
As illustrated in Figure 1 and 2(b), the freq. cross-self module in the decoder distinguishes between observed and unobserved frequency bins at the level of query construction. Specifically, for MHSA, the entire frequency axis is treated symmetrically, and all frequency bins participate as queries, keys, and values. In contrast, MHCA is applied only to the unobserved high-frequency region $\mathbf{Q}_{\text{ext}}$ when $f_{E}<f_{D}$, while the key–value are derived from the encoder features $\mathbf{Z}$ corresponding to the observed input bandwidth. Therefore, when $f_{E}=f_{D}$, no extension queries are instantiated and the MHCA path is bypassed entirely.
@@
Linformer (Wang et al., 2020) introduced linear projections of key-value to reduce the computations, while MLP-Mixer (Tolstikhin et al., 2021) entirely replaced MHSA with static linear operations. Motivated by these insights (Quan and Li, 2024), we incorporate a frequency linear projection (Figure 2(b)) to impose an inductive bias for the structural consistency of frequency bins on top of the benefits of dynamic attention. Since frequency bins exhibit consistent characteristics and considering the former studies (Wang et al., 2020; Quan and Li, 2024), we share the same projection layer across all modules and key-value mappings. Formally, for each head $h\hskip-2.84526pt=\hskip-2.84526pt1,\dots,\hskip-1.42262ptH$, given key-value $\mathbf{K}_{h,c},\hskip-1.42262pt\mathbf{V}_{h,c}\hskip-1.42262pt\in\hskip-1.42262pt\mathbb{R}^{T\times F}$ at channel $c$, a projection matrix $\mathbf{A}_{h}\hskip-2.27621pt\in\hskip-1.42262pt\mathbb{R}^{{F_{\text{max}}}\times F_{\text{proj}}}$ with dimension $F_{\text{proj}}$ and maximum bins $F_{\text{max}}$ is applied as
$\hskip-5.69054pt\tilde{\mathbf{K}}_{h,c}\hskip-2.84526pt$ $=$ $\hskip-2.84526pt[\mathbf{K}_{h,c},\mathbf{O}]\mathbf{A}_{h}\in\mathbb{R}^{T\times{F}_{\text{proj}}},\quad 1\leq c\leq C_{h},$ (1)
$\hskip-5.69054pt\tilde{\mathbf{V}}_{h,c}\hskip-2.84526pt$ $=$ $\hskip-2.84526pt[\mathbf{V}_{h,c},\mathbf{O}]\mathbf{A}_{h}\in\mathbb{R}^{T\times{F}_{\text{proj}}},\quad 1\leq c\leq C_{h},$ (2)
@@
The modular design of TF dual-path model further enables a seamless extension to streaming mode by replacing the time module with Mamba (Gu and Dao, 2024) blocks. Refer to Appendix B for detailed model configurations.